OpenStudy (anonymous):

Write the nth term of the following sequence in terms of the first term of the sequence. 10, 20, 40, . . . I'm almost positive this series is geometric because there is no solid difference. Is there a certain formula to obtaining the general term or does it vary for each one? Thanks y'all (:

OpenStudy (anonymous):

first tell which series

OpenStudy (anonymous):

isn't it geometric?

OpenStudy (anonymous):

because there is no difference of like adding

OpenStudy (anonymous):

ohh i meant common ratio

OpenStudy (anonymous):

Like, 10/20=1/2 & 20/40=1/2 ?

OpenStudy (anonymous):

nope..... do